# A Survey on Computational Solutions for Reconstructing Complete Objects by Reassembling Their Fractured Parts

### v1

@@ Abstract
Reconstructing a complete object from its parts is a fundamental problem in many scientific domains. The purpose of this article is to provide a systematic survey on this topic. The reassembly problem requires understanding the attributes of individual pieces and establishing matches between different pieces. Many approaches also model priors of the underlying complete object. Existing approaches are tightly connected problems of shape segmentation, shape matching, and learning shape priors. We provide existing algorithms in this context and emphasize their similarities and differences to general-purpose approaches. We also survey the trends from early non-deep learning approaches to more recent deep learning approaches. In addition to algorithms, this survey will also describe existing datasets, open-source software packages, and applications. To the best of our knowledge, this is the first comprehensive survey on this topic in computer graphics.
@@ 1 Introduction
Reconstructing a complete object from its parts is a fundamental problem in many scientific domains. In archaeology, we typically collect fragments of one or multiple objects at an archaeological site, and the goal is to restore the original objects. When people have bone fractures, we need to study how to assemble bone fragments to plan bone reduction surgery. This problem also arises in the 3D object/scene modeling, where we want to construct a complete object/scene from parts/objects. Other relevant problems are protein docking and human object interactions, in which the difference is that the transformation of each part becomes non-rigid. The purpose of this article is to provide a systematic survey on this topic, including algorithms, datasets, open source software packages, and applications. To the best of our knowledge, this is the first systematic survey on this topic.
@@
The reassembly problem is related to many shape analysis tasks. An important task is to segment the regions of the fractured surfaces and the original surfaces, which is a special shape segmentation problem. Another important task is to find correspondences between fractured regions of different pieces to find relative poses between them. This is related to scan matching, with the difference in that the correspondences are between complementary points. In addition, many approaches utilize priors of the underlying complete object, including symmetries and shape priors. This calls for approaches to analyze and encode such priors and use them to solve the reassembly problem.
@@
Similar to other problems in computer graphics, we have also witnessed the transitions from early non-learning based approaches to recent deep learning approaches. Early approaches have explored all aspects of the reassembly problem, including segmentation of fractured regions and original regions, matching fractured regions, and encoding priors (symmetries and shape spaces) of the underlying complete shape. However, they are restricted to hand-crafted features, so the entire systems do not generalize robustly. In addition, it is difficult to encode shape priors among objects with large geometrical and topological variations. These issues can be addressed using deep learning approaches. Besides the fact that the shape segmentation and shape matching problems become more robust, we also have approaches that perform shape segmentation and shape matching together, which benefit from end-to-end training in deep learning. In addition, with advances in learning 3D shape generative models, it is possible to learn shape priors from diverse shape collections to reassembly.
@@
Figure 1: Overview of this survey. (a) Single-piece processing and analysis. (b) Multi-piece analysis. (c) Template shape space and template-based reassembly. (d) Applications.
@@
Despite tight connections to relevant problems in shape analysis, applying existing approaches to the 3D reassembly problem is not straightforward. This is the case for both non-deep learning and deep learning based approaches. However, in the literature, people do not establish the reassembly as a unique sub-area in geometry processing and shape analysis. Our goal is to summarize state-of-the-art results in this domain systematically, seeking to establish the skeleton of this important sub-area.
@@ 1.1 Related Surveys
The most relevant survey is geometric analysis in cultural heritage [1], in which reassembling fractured objects is a central topic. The difference between our survey and [1] is two-fold. First, in contrast to their survey which focuses mostly on non-deep learning techniques, our survey covers recent advances using deep learning. Second, in contrast to the specific application of assembling fractured objects, this survey covers many applications, including bone fracture reduction, shape modeling by assembling parts, and applications in biology, geoscience, and paleontology.
Another relevant survey is [2], which studies how to decompose a shape into rigid parts. Its focus is different in the sense that it focuses on optimizing decompositions with respect to different geometric and physical criteria for manufacturing and fabrication. In contrast, we focus on solving the inverse problem of reconstructing complete objects from parts.
Our survey also discusses geometry processing techniques, including 3D reconstruction [3, 4], mesh segmentation [5, 6], primitive segmentation [7], shape descriptors [8] and correspondences [9], symmetry detection [10], and data-driven shape analysis and processing [11]. While the pipeline of reassembling parts uses these techniques, this article focuses on approaches tailored for this particular class of applications, which usually have modified inputs, outputs, and problem setups.
@@
There are also other relevant review articles. [15] surveys related approaches for each step in the pipeline that starts from reassembly to object completion, which also addresses missing regions. [16] present a survey article on the specific problem of reassembly of pottery fragments. [17] surveys related work on reassembly of defective cultural heritage objects. Our survey presents a comprehensive review of general fractured object reassembly and connections to relevant research results in geometry processing.
@@ 1.2 Paper Organization
We organize the reminder of this paper as follows. In Section 2, we provide a more detailed overview of this paper. In Section 3, we describe existing approaches for analyzing individual pieces. In Section 4, we discuss existing approaches for analyzing multiple pieces, focusing on problems of matching pairs of pieces and multiple pieces jointly. In Section 5, we introduce existing methods for encoding prior knowledge of the underlying complete object. In Section 6, we review approaches that utilize priors about the complete object to solve the reassembly problem. In Section 7, we discuss various applications of the reassembly problem. Section 8 present available datasets and software packages. In Section 9, we present current trends and future directions on this topic. Finally, we conclude this survey in Section 10.
@@ 2 Overview
This section presents an overview of this survey (See Figure 1).
Single piece processing and analysis (Sec. 3). Single piece processing and analysis concerns individual pieces of fractured objects. Topics include how to digitize physical fragments into digital 3D models and how to segment each fragment into regions of original surfaces and fractured surfaces. Although these topics are central topics of the field of geometry processing and shape analysis, we focus on techniques that are tailored for the particular characteristics of fragments of physical objects. We will cover both non-deep learning approaches and more recent deep learning approaches.
Multi-piece analysis (Sec. 4). Multi-piece analysis studies the critical problem of establishing correspondences between fragments to match them. We begin with pairwise matching techniques using geometric and textual features extracted from the fracture surfaces and original surfaces. We will survey global techniques that do not require prior knowledge of the relative pose between a pair of fragments and local techniques that start from reasonable initial relative poses. We then proceed to discuss methods that involve more than two pieces, in which there is a cycle-consistency constraint among the relative poses along a cycle of pieces. In addition to fragment matching, we will also discuss approaches that solve the correlated problems of fragment segmentation and fragment matching together. Similarly to single-piece analysis, we will cover both non-deep learning approaches and deep learning approaches.
@@
Applications (Sec. 7). After discussing technical approaches towards solving the reassembly problem, we will survey various applications of the reassembly problem. We will discuss five areas of fractured object reassembly application, including bone reduction, restoring fragments of thin-shells, walls, and building, and reassembly problems in science. Finally, we discuss the relevant problem of assembling parts to form complete objects and assembling objects to form 3D scenes.
Datasets and software packages (Sec. 8). We discuss publicly available datasets and open source software packages. Datasets include early small-scale datasets and recent large-scale real- and synthetic datasets.
Future directions (Sec. 9). The last paper of this survey summarizes existing approaches and discusses future directions. We will emphasize how to advance existing reassembly problems by developing better learning approaches, focusing on challenges in representations, data, and utilizing foundation models. Moreover, we will also discuss potential new applications of reassembly in natural and life sciences.
@@ 3.1 Reconstruction
Physical fragments must be captured and reconstructed into digital models before reassembling. This section discusses how fragmented pieces can be captured using various sensors, each requiring specific reconstruction techniques.
Laser scanning. As shown in Figure 2 (Top-Left), laser scanners capture the 3D geometry of fragments by emitting laser beams and measuring the time or phase difference of the reflected light. They allow for high-resolution geometry capture, achieving a physical distance of approximately $0.01$ mm between each captured point. The acquisition efficiency for capturing a piece, depending on the resolution selected, laser emission frequency, and the number of emitted laser lines, typically ranges from minutes to hours. For example, scanning a $10$ cm object using the low-frequency NextEngine scanner ($50K$ points per second) would take about an hour [18]. However, with high-frequency HandySCAN ($1.8M$ points per second), the process would only take a few minutes [19, 20]. Laser scanning is usually desirable for capturing high-quality geometry, particularly in outdoor environments. However, since it does not capture color information, other capturing modalities should be combined if chromatic data is needed. A significant drawback of these scanners is their high cost, though rental options are available. There are several representative datasets captured by laser scanning. Stanford University used laser scanners to collect the “Forma Urbis Romae” dataset as part of the Digital Michelangelo Project, capturing high-resolution 3D shapes of marble fragments for digital preservation and research [21]. Another significant dataset is generated through the “Reconstructing Thera Frescoes” project, where low-power laser scanning is employed to ensure accurate data capture while also protecting fragile frescoes from potential damage during the scanning process.[22]
Photogrammetry capturing. Color images captured by regular cameras can also be used to capture fragment geometry and texture; and recent studies have found that geometry reconstruction quality can be comparable to laser scanning, with a resolution of $0.15$  mm and an accuracy of $0.13$  mm [23, 24]. State-of-the-art multi-view stereo (MVS) algorithms reach pixel level accuracy, namely, $90\%$ reconstructed points have accuracy within about one pixel [25]. Therefore, the geometry quality can be further improved by capturing detailed images from a short range. As shown in Figure 2 (Top-Right), photogrammetry is well suited for field data collection because it does not require specialized skills to operate a scanner, is affordable and easy to set up. Designed for field collection without electricity, [26] proposed a rig to do photogrammetrical documentation of artifacts. Data collection for this protocol averaged approximately $20$ minutes, with a series of overlapping photographs (approximately $100$ per artifact) to be taken of the artifact from all angles [27]. To obtain even higher throughput, the batch method has been proposed by [24], achieving a throughput of about $730$ fragments per day. However, photogrammetry is sensitive to light and is not suitable for transparent material. The lighting needs to be closely controlled using two overhead fluorescent lamps or LEDs, which minimize the amount of shadows and regularized the amount of light across the surface[18]. Note that a list of “non-friendly” materials for photogrammetric purposes, such as water, glass, varnish, steel, and many more, can be found in [28]. In addition, selecting an appropriate focal length is essential to achieve optimal results in the digitalization process [29]. Wang et al. proposed a comprehensive dataset consisting of batches of 3D sherd scans, typically pottery fragments, accompanied by auxiliary information that aids in the alignment and reconstruction of the original artifacts through model registration techniques [24].
@@
Structured light capturing. As shown in Figure 2 (Bottom-Left), The main advantages of this technology are speed, making it suitable for applications where efficiency is crucial [32]. This method operates by projecting structured light onto an object and then analyzing the reflected light to determine the shape and dimensions of the object. The resolution of structured light scanners, which is around $0.1\sim 0.2$ mm, is slightly inferior to that of laser scanners. However, it can generally obtain better geometric detail and resolution in small details than digital photogrammetry. The operation time can range from 1 to 10 minutes, depending on the optical properties of the scanned object. In [33], 60 pottery fragments can be scanned per hour at the desired resolution. For some reflective pieces, such as flint, an experienced operator produced complete digital 3D images at the rate of $5\sim 6$ artifacts per hour [34]. Compared with laser, structured light can capture color information, but may struggle with shiny or dark surfaces. Therefore, they are typically well-suited for scanning small parts that require color and texture details, and for use in indoor environments. A representative project in this category is [35].
CT scanning. CT scanners are extremely useful in piece reconstruction because they allow 3D scanning not only of the outer surfaces of an object but also hidden internal structures [36, 37]. Unlike outer surface geometry, internal structures of the pieces are difficult to capture using the aforementioned scanning technology that relies on light reflection. The resolution is typically approximately $1\sim 500$ microns, a level of precision capable of capturing the fine topology [38]. To date, numerous studies have used CT to image rocks [39], stones [40], fossils [41], bone [42] and teeth [43]. Even though a high-resolution CT-scan can take several hours, the duration needed to scan a single part is manageable. [40] were able to scan up to 220 lithics at once in a relatively short period of time (ca. 2 hours) at a resolution of $140$ microns. Their tests suggested that this resolution is optimal for scanning many artifacts simultaneously while maintaining accuracy in the 3D shapes. The skull dataset [44] comprises high-resolution CT scans of both damaged and normal skulls, aiding in the study and reassembly of cranial anatomy. Another dataset [45] consists of high-resolution CT scans generated through helical scanning, providing detailed 3D reconstructions of anatomical structures stored in DICOM format.
@@
Registration and 3D reconstruction. In order to obtain the complete geometry of an object, it is necessary to scan from multiple angles and then use point cloud registration to align these scans in a common coordinate system. When reference markers are possible, such as small disc-shaped stickers [19] or unevenly spaced grid lines drawn on a turntable [18], they facilitate point cloud registration by serving as common reference points. When using such markers is infeasible, markerless registration such as Iterative Closest Point (ICP) algorithm and its variants are utilized to align the point clouds [33]. The quality of the initial setup in the Iterative Closest Point (ICP) algorithm is crucial, as it significantly influences the convergence rate and precision of the algorithm [46]. Therefore, careful manual annotations are essential for the optimum outcomes. In order to improve registration efficiency, [24] devised an automated and robust registration technique that leverages the bilateral boundary ICP, offering reliable initial relative pose estimates to initiate the ICP. After forming the dense point cloud, a mesh modeling algorithm, such as marching cubes [47] or Poisson reconstruction [48] is used to generate the surface, In contrast to the methods that record the scans before reconstructing the surface, [49] integrates both processes simultaneously, resulting in a noise-free and watertight solid model. Using joint registration, the method significantly improves reconstruction accuracy, ensuring better consistency when combining different views. The experimental results indicate that it can perfectly reconstruct the surfaces of archaeological fragments. This approach was used to reconstruct the fragments of the Octogan project [50].
@@
Batch processing in scanning improves efficiency and reduces the need for manual intervention. This approach is widely adopted in large-scale applications. [35] proposed a system to automatically scan multiple 3D objects using view and path planning algorithms, minimizing manual intervention. It enhances scanning efficiency and quality with structured light scanning and calibrated positioning, making it ideal for large-scale applications like archaeological documentation. [22] introduced a batch processing system that achieves efficient scanning through parallel operation of multiple devices, automated processing, and a fast matching algorithm. A single operator can manage up to four 3D scanners, while another uses a flatbed scanner to capture high-resolution images and normal maps. The system automatically aligns and processes data in the background, minimizing manual intervention. The fast geometry-based matching further optimizes efficiency, allowing the system to process approximately 10 fragments per hour. [24] presented another batch system to scan two sides of fragments and developed tailor registration and reconstruction techniques to process the paired scans.

### v2

@@ Abstract
Reconstructing a complete object from its parts is a fundamental problem in many scientific domains. The purpose of this article is to provide a systematic survey on this topic. This reassembly problem requires understanding the attributes of individual pieces and establishing matches between different pieces. Many approaches also model priors of the underlying complete object. Existing approaches are tightly connected problems of shape segmentation, shape matching, and learning shape priors. We provide existing algorithms in this context and emphasize their similarities and differences to general-purpose approaches. We also survey the trends from early procedural approaches to more recent deep learning approaches. In addition to algorithms, this survey will also describe existing datasets, open-source software packages, and applications. To the best of our knowledge, this is the first comprehensive survey on this topic in computer graphics.
@@ 1 Introduction
Reconstructing a complete object from its parts is a fundamental problem in many scientific domains. In archaeology, we collect fragments of one or multiple objects at an archaeological site, and the goal is to restore the original objects. In medical applications, we need to study how to assemble bone fragments to plan bone reduction surgery. This problem is also crucial in paleontology for reconstructing fossils and in forensics for evidence reconstruction. The challenge extends to 3D object/scene modeling, where we want to construct a complete object/scene from parts/objects. Other relevant problems are protein docking and human-object interactions, in which the difference is that the transformation of each part becomes non-rigid. The purpose of this article is to provide a systematic survey on this topic, including algorithms, datasets, open source packages, and applications. To the best of our knowledge, this is the first systematic survey on this topic.
@@
The reassembly problem is related to many shape analysis tasks. An important task is to segment and classify the parts’ surface into fractured and intact regions, which is a special shape segmentation problem. Another important task is to find correspondences between fractured regions of different pieces to find relative poses between them. This is related to scan matching, with the difference that the correspondences are between complementary points. In addition, many approaches utilize priors of the underlying complete object, including symmetries and shape priors. This calls for approaches to analyze and encode such priors and use them to solve the reassembly problem.
@@
Similar to other problems in computer graphics, we have also witnessed the transition from early procedural approaches to recent deep learning approaches. Early approaches have explored all aspects of the reassembly problem, including the segmentation of fractured regions and original regions, matching of fractured regions, and encoding of priors (symmetries and shape spaces) of the underlying complete shape. However, they are restricted to hand-crafted features, so entire systems do not generalize robustly. In addition, it is difficult to encode shape priors among objects with large geometrical and topological variations. These issues can be addressed using deep learning approaches. In addition to the fact that the shape segmentation and shape matching problems become more robust, we also have approaches that perform shape segmentation and shape matching together, which benefit from end-to-end training in deep learning. In addition, with advances in learning 3D shape generative models, it becomes possible to apply learned diverse shapes prior to reassembly.
@@
Figure 1: Overview of the methods presented in this survey. (a) Single-piece processing and analysis. (b) Multi-piece analysis. (c) Template shape space and template-based reassembly.
@@
Despite tight connections to relevant problems in shape analysis, applying existing approaches to the 3D reassembly problem is not straightforward. This is the case for both procedural and deep learning based approaches. While reassembly shares many technical components with other geometry processing tasks, it has not been formally recognized as a distinct sub-area within geometry processing and shape analysis. Through this survey, we aim to systematically organize state-of-the-art results in this domain and establish a structured foundation for this important research direction.
@@ 1.1 Related Surveys
The most relevant survey is geometric analysis in cultural heritage [1], in which reassembling fractured objects is a central topic. The difference between our survey and [1] is two-fold. First, in contrast to their survey which focuses mostly on procedural techniques, our survey covers recent advances using deep learning. Second, in contrast to the specific application of assembling fractured objects, this survey covers many applications, including bone fracture reduction, shape modeling by assembling parts, and applications in biology, geoscience, and paleontology.
Another relevant survey is [2], which studies how to decompose a shape into rigid parts. Its focus is different in the sense that it focuses on optimizing decompositions with respect to different geometric and physical criteria for manufacturing and fabrication. In contrast, we focus on solving the inverse problem of reconstructing complete objects from parts. [3] discussed techniques for reconstructing fractured archaeological textiles. It focuses more on using textiles as guidance for puzzle solving and archaeological reassembling in both 2D and 3D cases. Our work demonstrates more general cases and applications where textiles might not exist.
Our survey also discusses geometry processing techniques, including 3D reconstruction [4, 5], mesh segmentation [6, 7], primitive segmentation [8], shape descriptors [9] and correspondences [10], symmetry detection [11], and data-driven shape analysis and processing [12]. Although the process of reassembling parts uses these techniques, this article focuses on approaches tailored for this particular class of applications, which usually have modified inputs, outputs, and problem setups.
@@
There are also other relevant review articles. [16] survey related approaches for each step in the pipeline that starts from reassembly to object completion, which also addresses missing regions. [17] present a survey article on the specific problem of reassembly of pottery fragments. [18] survey related work on reassembly of defective cultural heritage objects. Our survey presents a comprehensive review of general fractured object reassembly and connections to relevant research results in geometry processing.
@@ 1.2 Paper Organization
We organize the remainder of this paper as follows. In Section 2, we provide a more detailed overview of this paper. In Section 3, we describe existing approaches for analyzing individual pieces. In Section 4, we discuss existing approaches for analyzing multiple pieces, focusing on problems of matching pairs of pieces and multiple pieces jointly. In Section 5, we introduce existing methods for encoding prior knowledge of the underlying complete object. In Section 6, we review approaches that utilize priors about the complete object to solve the reassembly problem. In Section 7, we discuss various applications of the reassembly problem. Section 8 presents available datasets and software packages. In Section 9, we present current trends and future directions on this topic. Finally, we conclude this survey in Section 10.
Figure 2: Overview of applications presented in this survey.
@@ 2 Overview
This section presents an overview of this survey (See Figure 1 and Figure 2).
Single piece processing and analysis (Sec. 3). Single piece processing and analysis concern individual pieces of fractured objects. Topics include how to digitize physical fragments into digital 3D models and how to segment each fragment into regions of original surfaces and fractured surfaces. Although these topics are central topics in the field of geometry processing and shape analysis, we focus on techniques that are tailored to the particular characteristics of fragments of physical objects. We will cover both procedural approaches and more recent deep learning approaches.
Multi-piece analysis (Sec. 4). Multi-piece analysis studies the critical problem of establishing correspondences between fragments to match them. We begin with pairwise matching techniques using geometric and textural features extracted from the fracture surfaces and original surfaces. We will survey global techniques that do not require prior knowledge of the relative pose between a pair of fragments and local techniques that start from reasonable initial relative poses. We then proceed to discuss methods that involve more than two pieces, in which there is a cycle-consistency constraint among the relative poses along a cycle of pieces. In addition to fragment matching, we will also discuss approaches that solve the correlated problems of fragment segmentation and fragment matching together. Similarly to single-piece analysis, we will cover both procedural approaches and deep learning approaches.
@@
Applications (Sec. 7). After discussing technical approaches to solving the reassembly problem, we will discuss various applications of the reassembly problem. We will discuss five areas of fractured object reassembly application, including bone reduction, restoring thin-shell, planar, and stone fragments, and reassembly problems in science. Finally, we discuss the relevant problem of assembling parts to form complete objects.
Datasets and open source (Sec. 8). We discuss publicly available datasets and open source packages. Datasets include early small-scale datasets and recent large-scale real- and synthetic datasets.
Future directions (Sec. 9). The last section of this survey summarizes existing approaches and discusses future directions. We will emphasize how to advance existing reassembly problems by developing better learning approaches, focusing on challenges in representations, data, and utilizing foundation models. Moreover, we will also discuss potential new applications of reassembly in natural and life sciences.
@@ 3.1 Reconstruction
Converting physical fragments into digital 3D models involves two key stages: digitization through various scanning technologies and registration to create complete 3D models. Although scanning choices influence data quality, the registration process shares fundamental algorithmic similarities with the assembly task itself. This section discusses how fragmented pieces can be captured using sensors, each of which requires specific reconstruction techniques.
Scanning Technologies. Several technologies are available for fragment digitization (Figure 3):
Laser Scanning captures high-resolution geometry ($\sim 0.01$ mm) by measuring reflected laser beams. High-frequency systems like HandySCAN ($1.8M$ points/second) enable rapid capture [19, 20], while lower-frequency options like NextEngine provide more affordable alternatives. Laser scanning excels in outdoor environments and has enabled major digitization efforts like Stanford’s “Forma Urbis Romae” dataset [21] and the “Reconstructing Thera Frescoes” project [22]. However, the lack of color capture and high equipment costs are notable limitations.
Photogrammetry achieves resolution comparable to that of laser scanning ($\sim 0.15$ mm) [23, 24] using a regular camera. State-of-the-art multi-view stereo algorithms achieve pixel-level accuracy for 90% of reconstructed points [25]. This approach is particularly suited for field work due to its low cost and minimal equipment requirements. Porter et al. [26] designed a specialized rig for field collection without electricity, requiring approximately 20 minutes and 100 photographs per artifact [27]. Although sensitive to lighting conditions and require controlled illumination [28], recent protocols have demonstrated a high throughput of 730 fragments per day [24]. However, photogrammetry struggles with certain materials such as water, glass, and steel [29], and requires careful selection of focal length [30].
Structured Light balances speed and accuracy ($\sim 0.1$-$0.2$ mm) by projecting patterns onto objects [31]. It can process 60 pottery fragments per hour [32] or 5-6 reflective artifacts per hour [33]. While capturing both geometry and color, it works best indoors and may struggle with shiny or dark surfaces. Notable applications include the work by Fan et al. [34].
CT Scanning is uniquely capable of capturing internal structures [35, 36] at micron-scale resolution ($1$-$500$ microns) [37]. It has been applied successfully to rocks [38], stones [39], fossils [40], bones [41], and teeth [42]. Recent work demonstrates batch processing of 220 artifacts in two hours at 140-micron resolution [39]. Important datasets include skull scans [43] and detailed anatomical structures [44].
@@
Registration and 3D reconstruction. In order to obtain the complete geometry of an object, it is necessary to scan from multiple angles and then use point cloud registration to align these scans in a common coordinate system. When reference markers are possible, such as small disc-shaped stickers [19] or unevenly spaced grid lines drawn on a turntable [28], they facilitate point cloud registration by serving as common reference points. When using such markers is infeasible, markerless registration such as the Iterative Closest Point (ICP) algorithm and its variants are utilized to align the point clouds [32]. A good initial alignment is essential for the ICP algorithm, as it directly affects how quickly and accurately the algorithm can find the final alignment [47]. Therefore, careful manual annotations are essential for optimum outcomes. In order to improve registration efficiency, [24] devised an automated and robust registration technique that leverages the bilateral boundary ICP, offering reliable initial relative pose estimates to initiate the ICP. After forming the dense point cloud, a mesh modeling algorithm, such as marching cubes [48] or Poisson reconstruction [49] is used to generate the surface. Unlike traditional methods that separate scanning and surface reconstruction, [50] combines these steps into a single process to create cleaner and more complete 3D models. Using joint registration, the method significantly improves reconstruction accuracy, ensuring better consistency when combining different views. The experimental results indicate that it can perfectly reconstruct the surfaces of archaeological fragments. This approach was used to reconstruct the fragments of the Octagon project [51].
@@
Batch processing in scanning improves efficiency and reduces the need for manual intervention. This approach is widely adopted in large-scale applications. [34] developed an automated system that can scan multiple 3D objects by intelligently planning camera views and scanning paths, reducing the need for human input. It enhances scanning efficiency and quality with structured light scanning and calibrated positioning, making it ideal for large-scale applications like archaeological documentation. [22] introduced a batch processing system that achieves efficient scanning through parallel operation of multiple devices, automated processing, and a fast matching algorithm. A single operator can manage up to four 3D scanners, while another uses a flatbed scanner to capture high-resolution images and normal maps. The system automatically aligns and processes data in the background, minimizing manual intervention. The fast geometry-based matching further optimizes efficiency, allowing the system to process approximately 10 fragments per hour. [24] presented another batch system to scan two sides of fragments and developed custom registration and reconstruction techniques to process the paired scans.